OpenStudy (anonymous):

an online lending service recently offered 48 month auto loans at 4.5% compounded monthly. if you can afford the payments of $521 how much can you borrow? what is the total interest?

OpenStudy (amistre64):

well, i have a formula ....

OpenStudy (anonymous):

that would help, thanks again! youre a life saver....

OpenStudy (amistre64):

\[Bk^n=P\frac{1-k^n}{1-k}\] solving for B is simple enough k is our compounding interest stuff (1+ i/12) and n is the number of compounding periods, 48

OpenStudy (preetha):

Welcome SNHU student. I am great fan of your school and your president, Paul Leblanc.

OpenStudy (anonymous):

Oh hi Preetha, it's a pretty good school... a little fast paced, but nice!

OpenStudy (amistre64):

how accurate does our amount need to be?

OpenStudy (anonymous):

two decimal places

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=521%281-%281%2B.045%2F12%29%5E%2848%29%29%2F%28%281%2B.045%2F12%29%5E%2848%29%281-%281%2B.045%2F12%29%29%29 well, if they expect you to use future value and present value tables, this isnt going to be ot their standards, but its more accurate

OpenStudy (amistre64):

the interest is just the total amount of payment, minus the amount of the loan. 48P = interest + principal

OpenStudy (anonymous):

oh :) thanks!

OpenStudy (amistre64):

the basic formula i use for most financial questions is: \[B_n=Bk^n-P\frac{1-k^n}{1-k}\] for n compounding periods, and some k compounding interest setup we can solve for payments, and B_o which is just B in the setup

OpenStudy (amistre64):

we can solve for n, and such as well ... its pretty handy

OpenStudy (anonymous):

do you have time to walk me through one ore question?

OpenStudy (anonymous):

if you buy a computer directly from the manufacturer for $3,099 and agree to repay it in 48 equal installments at 1.79% interest per month on the unpaid balance, how much are your monthly payments? how much total interest will be paid?

OpenStudy (preetha):

Don't forget to medal Amistre

OpenStudy (anonymous):

i believe i did

OpenStudy (preetha):

Thanks.

OpenStudy (anonymous):

yep! thank you guys!!!!